OpenStudy (anonymous):

x^2+(x+1)^2=61 solve equation

OpenStudy (anonymous):

expand on the left first

OpenStudy (anonymous):

i.e. "multiply out, combine like terms"

OpenStudy (anonymous):

x^2+(X+1)(x+1)-61=0

OpenStudy (anonymous):

keep going....

OpenStudy (anonymous):

X=-1, 61

OpenStudy (anonymous):

Im not sure what to do with the x^2

OpenStudy (anonymous):

expand what is \[(x+1)(x+1)=?\]

OpenStudy (anonymous):

you know, as the math teachers say "foil" i.e. multiply that stuff out do you know how to do it?

OpenStudy (anonymous):

x^2+2X+1

OpenStudy (anonymous):

sorry X*

OpenStudy (anonymous):

yeah so now we are at \[x^2+x^2+2x+1=61\] combine the like terms on the left

OpenStudy (anonymous):

then subtract \(61\) from both sides let me know when you get \[2x^2+2x-30=0\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oops \[2x^2+2x-60=0\]

OpenStudy (anonymous):

you there yet?

OpenStudy (anonymous):

ok got that

OpenStudy (anonymous):

divide by 2, get \[x^2+x-30=0\] then factor

OpenStudy (anonymous):

(x+6)(x-5)

OpenStudy (anonymous):

and X=-6 or 5

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw